OpenStudy (anonymous):

100 people are present in a party. If each of them must do a handshake with all the other people, how many handshakes should be done? Please explain how to get the answer, thanks!

OpenStudy (anonymous):

you need 2 people for an hand shake and u have 100 people so i think that its C(100,2) but i am not sue of it..

OpenStudy (anonymous):

\[\sum_{n=1}^{99}n=\frac{1}{2}\times99\times100=4950\]

OpenStudy (anonymous):

Could you solve it using factorials, please?

OpenStudy (anonymous):

hmm i found the answer as 9900

OpenStudy (amistre64):

if there are 3 people: 1,2 ; 1,3 ; 2,3 .... is that right?

OpenStudy (anonymous):

Think about it like this: the first person has 99 people to shake hands with, the second person has 98 because they've already shaken hands with the first, the third has 97...

OpenStudy (anonymous):

yea but its harder to calculate this like that... so i think C(100,2) better way of solve

OpenStudy (amistre64):

5054 if we do that ..... add all the numbers from 1 to 100 :)

OpenStudy (anonymous):

No, all the numbers from 1 to 99.

OpenStudy (amistre64):

n(n+1) ------ :) 2

OpenStudy (amistre64):

ack .... 99 then lol

OpenStudy (amistre64):

wouldnt the last guy have noone to shake hands with?

OpenStudy (anonymous):

mhmh yea but there is 100 people and they and we need 2 people to shake hands so that it should be 100 x 99 /2,

OpenStudy (anonymous):

Yeah, the last guy has no-one to shake hands with, that's why it's sum of the first 99 integers, not the first 100.

OpenStudy (anonymous):

yea you right

OpenStudy (amistre64):

5.4321 4.321 3.21 2.1 1.0 i get 5 people with 10 shakes 5(4)/2 = 10 8.7654321 7.654321 6.54321 5.4321 4.321 3.21 2.1 1.0 8 people with 28 shakes 8(7)/2 = 28 I spose it works :) 100(99)/2 = 99*50 = 4950 yay!!

OpenStudy (anonymous):

Glad we agree.

OpenStudy (amistre64):

factorial would be: 100! 100*99 ---------- = ------- 2! (100-2)! 2

OpenStudy (anonymous):

its the same with C(100,2)